OpenStudy (migitmack):

How can I find the area of rectangle ABCD if I know the perimeter and the height.(The perimeter is 54 cm). The height is 15cm.

OpenStudy (anonymous):

Well, you know the height right?

OpenStudy (migitmack):

yea

OpenStudy (anonymous):

there are two sides, for the 'height' so... 54-15-15=24

OpenStudy (anonymous):

how would you use 24 to get the answer to the last two sides?

OpenStudy (migitmack):

divide it by 54?

OpenStudy (anonymous):

There are two sides left, and 24 left. Divide it by 2 :)

OpenStudy (migitmack):

oh XD kk

OpenStudy (migitmack):

i knew i had to divide though

OpenStudy (anonymous):

You did :)

OpenStudy (migitmack):

24/2=12

OpenStudy (migitmack):

so that means 12*15=180 which is the AREA! :D

OpenStudy (anonymous):

Great! Now you have the height and the width. The area is cake after this :D

OpenStudy (anonymous):

Yeah! :D

OpenStudy (migitmack):

YEAH!! ty

OpenStudy (anonymous):

<3 of course